OpenStudy (anonymous):

y= 3/4 x - 4 and i need to create a table when x = 0,1,2,3 what does y equal

OpenStudy (jdoe0001):

\(\large y = \cfrac{3}{4}{\color{red}{ x}}-4\qquad \qquad \begin{array}{cccccl} x&\qquad &y\\ \hline\\ 1&&\frac{3}{4}\cdot {\color{red}{ 1}}-4\\ 2&&\frac{3}{4}\cdot {\color{red}{ 2}}-4\\ 3&&\frac{3}{4}\cdot {\color{red}{ 3}}-4\\ 4&&\frac{3}{4}\cdot {\color{red}{ 4}}-4 \end{array}\)

OpenStudy (jdoe0001):

hmm oh.... skipped 0... but anyow... there, just plug in 0 and get "y"